OpenStudy (anonymous):

Use a linear approximation (or differentials) to estimate the given number. (Use the linearization of 1/x) 1/99

OpenStudy (anonymous):

everything deleted!

OpenStudy (anonymous):

That's something wrong

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Let me redo it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

L(100)=F(100)+F'(100)(-1) 0.0101

OpenStudy (anonymous):

Let F(x)=1/x, we know F(100) but we want to know F(99) so we use linearization L(x)=F(x)+F'(x)dx 1/x - 1/x^2 (-1) dx is -1 because we are going from 100 to 99

OpenStudy (anonymous):

so what is x

OpenStudy (anonymous):

99 is x?